# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v1

@@ 6.2. Future Work
Model: Extending the IAI Model. The IAI model intentionally abstracts at a high level to capture the interplay between text prompts and interactions. This abstraction aims for generality, but extensions can enrich the model for more fine-grained analysis of specific design questions. At the entity level, entities can be refined or expanded to capture richer system dynamics. For instance, as noted in Sec. 3.1, while context is currently subsumed under the artifact entity, making it explicit would support the analysis of how retrieval-augmented systems use user provenance. In addition, the “human” entity could generalize into an actor role, instantiated by either humans or AI agents, enabling representation of mixed-initiative or multi-AI agent workflows (Subramonyam et al., 2024; Li et al., 2024). At the level of relations, finer distinctions can also sharpen analysis. For example, the link from interaction to augmented instruction manifests differently across tools, ranging from sliders and widgets to sketching or structured graph editing. Making such variations explicit would not only capture existing diversity but also open design opportunities, such as composable augmented instruction libraries for domain-specific workflows (Setlur et al., 2020), adaptive timing of system responses (Gajos et al., 2006), or model-assisted widget generation (Chen et al., 2025b).

### v2

@@ 6.2. Future Work
Model: Extending the IAI Model. The IAI model intentionally abstracts at a high level to capture the interplay between text prompts and interactions. This abstraction aims for generality, but extensions can enrich the model for more fine-grained analysis of specific design questions. At the entity level, entities can be refined or expanded to capture richer system dynamics. For instance, as noted in Sec. 3.1, while context is currently subsumed under the artifact entity, making it explicit would support the analysis of how retrieval-augmented systems use user provenance. In addition, the “human” entity could generalize into an actor role, instantiated by either humans or AI agents, enabling representation of mixed-initiative or multi-AI agent workflows (Subramonyam et al., 2024; Li et al., 2024). At the level of relations, finer distinctions can also sharpen analysis. For example, the link from interaction to augmented instruction manifests differently across tools, ranging from sliders and widgets to sketching or structured graph editing. Making such variations explicit would not only capture existing diversity but also open design opportunities, such as composable augmented instruction libraries for domain-specific workflows (Setlur et al., 2020), adaptive timing of system responses (Gajos et al., 2006), or model-assisted widget generation (Chen et al., 2025a).
@@
Limitations. Our contributions should be interpreted within several boundaries. First, the IAI model and the analyzed atomic paradigms currently focus on single-user, single-agent interactions and do not yet formalize coordination among multiple GenAI nodes or the construction of shared augmented instructions. Second, the taxonomy is derived from a UI-centric corpus of 66 tools, which, although diverse, cannot fully represent the rapidly evolving GenAI ecosystem. Third, validation relies on qualitative synthesis and illustrative case studies rather than empirical user or designer studies. Future work could include controlled evaluations and real-world deployments to assess practical adoption and impact.
@@ 7. Conclusion
Acknowledgements.
This research is part of the AFMR collaboration supported by Microsoft Research. This work is partially supported by Hong Kong RGC GRF grant (16218724).